OpenStudy (anonymous):

Find the approximate value of the circumference of a circle with the given radius. Use = 3.14. Round your answer to the nearest tenth. 4 inches

OpenStudy (anonymous):

The radius is 4 inches?

OpenStudy (anonymous):

Well, the Circumference formula is C = pi * d pi is 3.14 d is the diameter if the radius is 4, the diameter is 8 because the diameter is twice the size of the radius. So your answer will be: C = 3.14 * 8 C = 25.12

OpenStudy (anonymous):

After you round it to the nearest tenth it's 25.1